1.Find two consecutive integers such that the sum of one-fourth of the first and one-fifth of the second is 11.
2. The denominator of a certain fraction is 6 more than the numerator, and the fraction is equivalent to one-third. Find the numerator.

Answer by mananth(16946)

let the integers be x & x+1
x/4 = (x+1)/5 =11
multiply equation by 20
5x+4(x+1)=220
5x+4x+4 =220
9x+4=220
9x=216
x=24
24,25
------------
x/(x+6) =1/3
cross multiply
3x=x+6
3x-x=6
2x=6
x=3
numerator =3
denominator = 3+6 =9
fraction 3/9